# Effects of the Cyber Resilience Act (CRA) on Industrial Equipment Manufacturing Companies

### v1

@@ Abstract
The Cyber Resilience Act (CRA) is a new European Union (EU) regulation aimed at enhancing the security of digital products and services by ensuring they meet stringent cybersecurity requirements. This paper investigates the challenges that industrial equipment manufacturing companies anticipate while preparing for compliance with CRA through a comprehensive survey. Key findings highlight significant hurdles such as implementing secure development lifecycle practices, managing vulnerability notifications within strict timelines, and addressing gaps in cybersecurity expertise. This study provides insights into these specific challenges and offers targeted recommendations on key focus areas, such as tooling improvements, to aid industrial equipment manufacturers in their preparation for CRA compliance.
Keywords:
CRA, Regulation, Challenges, Survey, IEC-62443, Standards
@@ 1 Introduction
The EU Cyber Resilience Act (CRA) [8] sets common security requirements for all products with digital elements within the EU. This broad scope intentionally covers all software and hardware products, as well as their remote data processing solutions, with exceptions for medical devices, aeronautical equipment, etc. which are already subject to similar existing EU regulations. The primary goal of the CRA is to enhance cybersecurity across the European Union by ensuring that products with digital elements are designed, developed, and maintained with robust security measures. While the act imposes requirements on many stakeholders, a significant portion of these requirements are directed towards manufacturers. The requirements for manufacturers can broadly be classified into software development lifecycle (SDL) process requirements, technical requirements, vulnerability management and notification requirements, and preparation of technical and user documentation.
@@
Figure 1: CRA timeline
Given the substantial requirements imposed on manufacturers and the relatively short timeline for compliance, we conducted a quick survey to pinpoint the specific challenges faced by industrial equipment manufacturing companies. Our goal was to identify their pain points and offer targeted recommendations to aid in their preparation for compliance.
@@ 2 Previous Work on CRA Challenges
Szedlak et al. [15] examine the awareness and preparedness of SMEs for CRA. Their findings reveal significant disparities in CRA awareness and readiness, with only 12.3% of SMEs being aware of the CRA compared to 83.5% of very large enterprises.
@@
While previous work highlighted several challenges associated with CRA, our survey focuses on investigating the specific challenges that CRA poses for industrial equipment manufacturing companies. By identifying these challenges, we aim to provide a comprehensive understanding of how CRA impacts industrial equipment manufacturers and offer insights into effective strategies in key focus areas for seamless CRA compliance.
@@ 3 CRA and IEC 62443
In the survey, our questions to participants primarily focused on the IEC 62443-4-1 SDL process requirements necessary for CRA compliance, since following this process ensures that a suitable risk analysis is performed and that technical requirements, such as confidentiality and integrity protection of transmitted and stored data, are met. In addition to SDL process challenges, the survey included questions about the challenges of vulnerability notification and user instructions documentation requirements.
@@ 4 Study Design
Between June and August 2024, we conducted a survey to gather insights into the compliance challenges faced by industrial equipment manufacturing companies. The survey utilized an online questionnaire with 44 questions focused on CRA and IEC 62443-4-1 compliance. The questionnaire included both multiple-choice and freeform questions. Participants completed the questionnaire during one hour video meetings, where they could provide feedback using the think-aloud protocol. This approach helped capture comments that might otherwise have been missed, especially for the freeform qualitative questions. All meetings were hosted by the same author, and the survey questionnaire was shared with participants prior to the video meeting. Each session concluded with a discussion to gather broader feedback on challenges and potential solutions.
@@
12345123450123456IEC 62443-4-1FamiliarityCRAFamiliarityFamiliarity:1. Not familiar at all2. Slightly familiar3. Moderately familiar4. Quite familiar5. Very familiar 3333333366661111222255554444Familiarity levelAnswer count0123456Answer count
@@ 5 Challenges
Practice 1Practice 2Practice 3Practice 4Practice 5Practice 6Practice 7Practice 8Technical requirementsSW Dev LifecycleVuln. ManagementVuln. NotificationConformity assessmentUser manuals012345678IEC 62443-4-1ChallengesCRAChallenges77772222777733334444333366663333222266665555666666663333IEC 62443-4-1 and CRA componentsAnswer count012345678Answer count
@@
The CRA requirement text “Products with digital elements shall be designed, developed, and produced in such a way that they ensure an appropriate level of cybersecurity based on the risks” implies the need for a mature secure software development lifecycle (SDL). Many survey participants found this challenging, echoing the difficulties observed with IEC 62443-4-1 requirements discussed earlier. Establishing and enforcing an SDL across an organization is a substantial undertaking. Vulnerability notification requirements were also highlighted as a significant challenge, as many participants had little prior experience with such practices. The CRA deadline of notifying within 24 hours of an actively exploited vulnerability was perceived as stringent. Some participants even expressed doubts about their ability to detect actively exploited vulnerabilities promptly. Similarly, participants were uncertain whether their vulnerability management processes and tools were sufficiently mature to meet CRA expectations. Conformity assessment was another major area of concern. Participants were unsure whether their products required external assessment by a notified body or if internal self-declaration would suffice based on their product type. Even when self-declaration was deemed acceptable, participants questioned whether their quality teams, which currently oversee self-declaration and CE labeling, were adequately equipped to evaluate compliance with the cybersecurity requirements of the CRA. Finally, the pattern observed earlier repeated, the technical requirements of CRA, such as confidentiality and integrity protection of stored and transmitted data, were not generally perceived as challenging.
@@
Cost Impact and Organizational Changes When asked about the estimated cost impacts of CRA compliance, as evident in figure 4, 50% of participants expect costs to remain under 20%, 25% anticipate costs to increase by 20–40%, and, surprisingly, 25% foresee no cost impact at all. Similarly, when participants were asked about potential changes in employee roles, also shown in figure 4, the majority (50%) expect roles to remain unchanged after the CRA comes into force, while 33.3% anticipate role changes, and 16.7% remain uncertain. These findings suggest that CRA may not significantly impact employee roles or organizational costs. This could indicate that manufacturers are either reasonably prepared or have resigned themselves to achieving compliance using their existing resources without expecting additional budget or support for structural changes.
[…]
Role ChangesNo ImpactUnder 20%20-40%Over 40%012345633336666333300Cost ImpactAnswer count0123456Answer count
@@ 6 Tackling CRA Challenges
As is evident from the survey results, industrial equipment manufacturers anticipate several challenges for CRA compliance. In this section, we focus on three key areas, and provide practical recommendations for tackling the challenges.
@@
The importance of vulnerability management tools was a recurring theme among participants. Many recognized the value of standard SBOM formats like CycloneDX [2] and tools for tracking vulnerabilities based on SBOM data. While vulnerability management tools are necessary and readily available, companies should also prepare for centrally aggregating vulnerabilities from multiple sources and constructing a hierarchy of vulnerabilities that accurately reflects the composition of their product components. Adopting the Common Vulnerability Scoring System (CVSS) for impact assessment is essential too. To reduce noise from false positives or low-impact vulnerabilities, organizations should consider advanced risk-rating methods, including AI-driven analysis if necessary.
@@ 7 Conclusion
We conducted a survey of industrial equipment manufacturing companies to explore the challenges they anticipate as they prepare for CRA compliance. The survey revealed several key concerns, including the tight timelines for vulnerability notifications, the requirement to support timely automatic security updates with opt-out options, the need for updated user manuals, and the complexities associated with using open-source software. Through freeform discussions during the survey, participants shared ideas for addressing these challenges, which we have synthesized into practical recommendations across three critical areas: improving communication within organizations and among stakeholders, addressing the shortage of cybersecurity experts, and enhancing security tooling. These findings provide valuable insights that can help other companies assess whether their challenges align with those identified by their peers. This knowledge, along with the practical recommendations, should support industrial equipment manufacturing companies, as well as other organizations, in proactively preparing for CRA compliance.

### v2

@@ Abstract
The Cyber Resilience Act (CRA) is a new European Union (EU) regulation aimed at enhancing the security of digital products and services by requiring them to meet stringent cybersecurity requirements. To understand the practical implications of CRA for industrial equipment manufacturing companies, a survey was conducted to identify key challenges. The results revealed significant hurdles, including the implementation of secure development lifecycle practices, managing vulnerability notifications within strict timelines, and addressing gaps in cybersecurity expertise. Based on these findings, the paper offers targeted recommendations in key focus areas such as vulnerability management and tooling improvements to support industrial equipment manufacturers in preparing for CRA compliance. [^1] This is the accepted manuscript version. The final published version is available at https://doi.org/10.1007/978-3-032-00630-1_12
Accepted Manuscript
This version of the article has been accepted for publication in ARES 2025 workshop. The final version is available online at SpringerLink.
@@ 1 Introduction
Motivated by the growing number of cyberattacks targeting connected devices and the lack of consistent baseline cybersecurity practices among manufacturers, the European Union (EU) Cyber Resilience Act (CRA) [4] sets common security requirements for all products with digital elements within the EU. This broad scope intentionally covers all software and hardware products, as well as their remote data processing solutions, with exceptions for medical devices, aeronautical equipment, etc. which are already subject to similar existing EU regulations. The primary goal of CRA is to enhance cybersecurity across the European Union by ensuring that products with digital elements are designed, developed, and maintained with robust security measures. While the act imposes requirements on many stakeholders, a significant portion of these requirements are directed towards manufacturers. The requirements for manufacturers can broadly be classified into software development lifecycle (SDL) process requirements, technical requirements, vulnerability management and notification requirements, and preparation of technical and user documentation.
@@
Non-compliance with the CRA can have serious consequences for manufacturers, including restricted market access, mandatory product recalls, or financial penalties. Given the extensive regulatory obligations and the relatively short timeline for full compliance, we conducted a survey to identify the specific challenges faced by industrial equipment manufacturing companies. These companies typically have well-established engineering practices, large development organizations, and structured product lifecycle processes already in place. However, their products often include deeply embedded digital elements integrated into complex, safety-critical, bespoke system installations, such as cranes and mining equipment. In such contexts, supporting requirements like automatic security updates and configuration changes are non-trivial. Additionally, organizational inertia and the need to coordinate changes across multiple product lines, regulatory environments, and customer expectations can make the adoption of CRA particularly challenging. Our goal was to identify concrete pain points and offer targeted recommendations to aid in their preparation for compliance.
The rest of this paper is organized as follows. Section 2 reviews prior work on the challenges associated with CRA implementation. Section 3 examines the IEC 62443 series of standards and discusses how they align with CRA requirements. The study design, including participant background and familiarity, is described in Section 4. Section 5 presents the key challenges identified by the surveyed industrial equipment manufacturers. In Section 6, we provide targeted recommendations to address the challenges identified. Finally, Section 7 concludes the paper and summarizes the main findings.
@@ 2 Previous Work on CRA Challenges
Szedlak et al. [19] examine the awareness and preparedness of Small and Medium-sized Enterprises (SMEs) for CRA. Their findings reveal significant disparities in CRA awareness and readiness, with only 12.3% of SMEs being aware of the CRA compared to 83.5% of very large enterprises.
@@
While previous work highlighted several challenges associated with CRA, our survey focuses on investigating the specific challenges that CRA poses for industrial equipment manufacturing companies. By identifying these challenges, we aim to provide an understanding of how CRA impacts industrial equipment manufacturers and offer insights into effective strategies in key focus areas for seamless CRA compliance.
@@ 3 CRA and IEC 62443
In our survey, a major component of the questions focused on the IEC 62443-4-1 secure development lifecycle (SDL) process requirements relevant to CRA compliance. Organizations that have already adopted IEC 62443 practices are generally well positioned to meet many of the CRA’s expectations, especially in areas related to secure development, risk-based design, and technical protections for the confidentiality and integrity of transmitted and stored data. Many of the participating companies were indeed applying IEC 62443 development practices internally. However, CRA introduces additional requirements not fully covered by the current IEC 62443 series. These include fixed timelines for vulnerability notification, mandatory public disclosure under specific conditions, and more detailed user documentation expectations. Therefore, in addition to questions on SDL process challenges, our survey also explored participants’ views on meeting CRA’s vulnerability notification obligations and user documentation requirements, which extend beyond what is explicitly addressed in IEC 62443. To deliver CRA-compliant products by December 2027, manufacturers will need to develop and integrate these additional capabilities into their existing processes.
@@ 4 Study Design
Between June and August 2024, we conducted a survey to gather insights into the compliance challenges faced by industrial equipment manufacturing companies. The survey utilized a detailed online questionnaire comprising 44 questions that covered various aspects of CRA and IEC 62443-4-1 compliance. The questionnaire included both multiple-choice and freeform questions, focusing on topics such as the role of EU standards within the organization, familiarity with IEC 62443 and CRA, existing secure software development lifecycle (SDL) practices, whether companies already follow IEC 62443-4-1, perceived difficulty in implementing specific practices, challenges in meeting CRA requirements such as vulnerability management and notification, expected role and budget changes, and anticipated compliance timelines.
Participants completed the questionnaire during one-hour video meetings, where they could provide additional feedback using the think-aloud protocol. This approach helped capture verbal comments that might otherwise have been missed, especially in response to the freeform qualitative questions. All meetings were hosted by the same author, and the questionnaire was shared with participants prior to the video call. Each session concluded with an open-ended discussion to gather broader feedback on challenges and potential solutions. Participants were free to refer to the IEC 62443 standard and the CRA text while answering the questions. If they asked for clarification about the meaning of specific requirements or practices, the facilitator provided contextual explanations and examples to ensure a shared understanding.
@@
The survey responses were analyzed using a qualitative coding approach commonly used in exploratory user studies. One of the authors reviewed the meeting recordings and annotated relevant open-ended comments that were not captured in the survey form. These responses were then coded thematically to identify recurring topics and pain points. Both text-based and spoken responses were included in the frequency analysis to determine how often specific concerns or challenges were mentioned across participants. Observed patterns and consistent themes were synthesized to form the basis for the practical recommendations presented later in the paper.
@@ 5 Challenges
The CRA requirement text “Products with digital elements shall be designed, developed, and produced in such a way that they ensure an appropriate level of cybersecurity based on the risks” implies the need for a mature secure software development lifecycle (SDL). Many survey participants found this challenging, echoing the difficulties observed with the IEC 62443-4-1 requirements discussed earlier. Establishing and enforcing an SDL across an organization is a substantial undertaking.
Participants were also uncertain whether their vulnerability management processes and tools were sufficiently mature to meet CRA expectations. Several respondents highlighted challenges such as generating accurate and up-to-date SBOMs, running and integrating appropriate vulnerability scanning tools, and effectively triaging and prioritizing vulnerabilities. Even when vulnerabilities are identified, issuing timely patches can be difficult in practice, particularly in industrial settings where the update cycles are inherently slower due to end-user demands for high availability and minimal downtime. These operational realities were identified as barriers to fulfilling CRA’s requirements for timely remediation and clear vulnerability reporting.
Similarly, vulnerability notification requirements were also highlighted as a significant challenge, as many participants had little prior experience with such practices. The CRA deadline of notifying within 24 hours of an actively exploited vulnerability was perceived as stringent. Some participants even expressed doubts about their ability to detect actively exploited vulnerabilities promptly.
Conformity assessment was another major area of concern. Participants were unsure whether their products required external assessment by a notified body or if internal self-declaration would suffice based on their product type. Even when self-declaration was deemed acceptable, participants questioned whether their quality teams, which currently oversee self-declaration and CE marking, were adequately equipped to evaluate compliance with the cybersecurity requirements of the CRA. In addition, it was not clear to participants whether sector-specific standards based on IEC 62443 would be sufficient for demonstrating conformity under the CRA. For example, the maritime sector has adopted the International Association of Classification Societies (IACS) Unified Requirements UR E26 [9] and UR E27 [10], which apply to new ships contracted from 1 January 2024 and draw on IEC 62443. However, it was unclear to participants whether compliance with or certification against these standards will fully meet all CRA requirements or be sufficient for self-declaration or assessment by a notified body.
@@
Cost Impact and Organizational Changes When asked about the estimated cost impacts of CRA compliance, as shown in Figure 4, 50% of participants expect costs to remain under 20%, 25% anticipate costs to increase by 20–40%, and, surprisingly, 25% foresee no cost impact at all. The question asked participants to estimate how much additional cost CRA compliance would introduce relative to their current product development budgets, including expenses related to personnel, tooling, and process adjustments. Similarly, when participants were asked about potential changes in employee roles, also shown in figure 4, the majority (50%) expect roles to remain unchanged after the CRA comes into force, while 33.3% anticipate role changes, and 16.7% remain uncertain. These findings suggest that CRA may not significantly impact employee roles or organizational costs. This could indicate that manufacturers are either reasonably prepared or have resigned themselves to achieving compliance using their existing resources without expecting additional budget or support for structural changes.
@@ 6 Tackling CRA Challenges
As is evident from the survey results, industrial equipment manufacturers anticipate several challenges in preparing for CRA compliance. This section focuses on the main areas where participants discussed practical difficulties and explores recommendations that emerged during the survey conversations to help address these challenges.
@@
The importance of vulnerability management tools was a recurring theme among participants. Many recognized the value of standard SBOM formats like CycloneDX [2] and tools for tracking vulnerabilities based on SBOM data. While vulnerability management tools are necessary and readily available, companies should also prepare for centrally aggregating vulnerabilities from multiple sources and constructing a hierarchy of vulnerabilities that accurately reflects the composition of their product components. Adopting a vulnerability risk scoring methodology grounded in Common Vulnerability Scoring System (CVSS) is essential too. To reduce noise from false positives or low-impact vulnerabilities, organizations should consider advanced risk-rating methods, including AI-driven analysis if necessary.
@@
Vulnerability Management and Notification Practices In addition to tooling, manufacturers should develop and document internal vulnerability management processes that clearly define roles, responsibilities, and workflows from detection to resolution. These should cover vulnerability identification using scanning tools, risk triage to assess severity and impact, and escalation procedures for critical or high-priority findings. The process must also include handling false positives, assigning remediation tasks, and coordinating actions across relevant teams such as product owners, security engineers, and operations. It should define criteria for when public vulnerability notification is required under CRA, and assign roles to ensure timely and accurate reporting to authorities such as ENISA or national CSIRTs. The process should also account for compensating controls when immediate fixes are not feasible, and include steps for planning, approving, and deploying fixes through the organization’s existing change management procedures. It must also support verification of fixes, handle residual risk through formal exception workflows, and be overseen by a central governance structure that regularly reviews critical findings and notification decisions. Where available, manufacturers are encouraged to base their process on harmonized standards that are intended to demonstrate CRA compliance within their sector. In the absence of such a standard, ISO/IEC 29147 [11] and ISO/IEC 30111 [13] can serve as useful inputs for designing an effective internal process.
@@ 7 Conclusion
We conducted a survey of industrial equipment manufacturing companies to explore the challenges they anticipate as they prepare for CRA compliance. The survey revealed several key concerns, including the tight timelines for vulnerability notifications, the requirement to support timely automatic security updates with opt-out options, the need for updated user manuals, the complexities associated with using open-source software, and uncertainty around conformity assessment procedures and the applicability of sector-specific standards. Through freeform discussions during the survey, participants shared ideas for addressing these challenges, which we have synthesized into practical recommendations across several focus areas. These include improving internal vulnerability management processes and notification readiness, enhancing communication between cybersecurity, quality, and leadership teams, addressing gaps in cybersecurity expertise through training and shift-left strategies, integrating supportive tooling into development workflows, and ensuring that conformity assessment processes are well understood and documented. While organizations already following IEC 62443 are better prepared, CRA introduces additional obligations, particularly in the areas of documentation, vulnerability notification, and timely update readiness, that must be addressed now to ensure compliance by December 2027. The findings and recommendations presented in this paper offer practical guidance to help industrial equipment manufacturers and other organizations facing similar challenges prepare effectively for CRA compliance.